OpenStudy (anonymous):

MULTIPLE CHOICE!! In a population of plants, the allele for long stems (S) is completely dominant over the allele for short stems (s). If 35% of the population has short stems, calculate the percentage of the population that is expected to be heterozygous (Ss). 48 percent 59 percent 41 percent 17 percent

OpenStudy (anonymous):

41 percent

OpenStudy (rogue):

I think its 48% \[p + q = 1\]\[p^2 + 2pq + q^2 = 1\]\[q^2 = 0.35 \rightarrow q = 0.59 \rightarrow p = 0.41\]\[Heterozygous: 2pq = 2 \times 0.41 \times 0.59 = 0.48 \rightarrow 48%\]